OpenStudy (anonymous):

Suppose (1, -5/2) is the midpoint of the line segment AB and suppose A is the point (5,5). Find B.

OpenStudy (nenadmatematika):

\[(5+x)\div2=1 x =-3 (5+y)/2=-5/2 y=-10 so B (-3,-10)